# InfiniteHBD: Building Datacenter-Scale High-Bandwidth Domain for LLM with Optical Circuit Switching Transceivers

### v3

@@ Appendix B GPT-MoE Architecture
Appendix C Theoretical analysis of wasted GPU ratio for InfiniteHBD
The count of backup lines as $22K-$ 2 will significantly influence the fault tolerance of InfiniteHBD. We use the expectation of waste ratio caused by GPU failure and fragmentation problem to evaluate this design, the result is shown in Table 7.
For one single working node in the middle of line, the count of breakpoints $B$ on its two sides has the expectation as:
@@ Appendix C Theoretical analysis of wasted GPU ratio for InfiniteHBD
Where $P_{s}$ is the fail probability of GPU node, and $\eta$ is count of nodes. The expectation of breakpoints count is:
@@
When the distance between two nodes is $\beta\geq K$, the breakpoints among them can be calculated as independent. Once the distance $\beta<K$, as all nodes in this range are connected to these two nodes, there will be no breakpoints between them. So, the expectation is less than two independent nodes. Then,
@@
For a LLM job which require a ring communication size (TP .etc) as $N_{t}$, InfiniteHBD will cut the whole line topology into several sub lines with the length of $N_{t}/R$. Once InfiniteHBD is cutting a new sub line from the remaining nodes in the line, all $N_{t}$ GPU will be wasted when one break point exist in the middle of this sub line required, shown in Figure 15. Then the expectation for waste GPU caused by one single break point is:
@@
In our trace for a 348 days long pre-train job on 10K-GPU, the p99 failure rate for 8-GPU nodes is 7.92%. If a TP32 jobs is running on InfiniteHBD, we can get the upper bond for waste ratio expectation for various configuration in Table 7.
@@
R=4 $9.22\%$ $0.37\%$ $1.52\times 10^{-4}$
R=8 $30.11\%$ $2.38\%$ $0.19\%$
Table 7. Upper bond for waste ratio expectation of GPU, where GPU failure rate is 1.03% and $N_{t}$ is 32
As shown in the table, for 4-GPU node ($R=4$) 3 bundles ($K=3$) design, the additional waste of GPU is less than 0.37%, while the waste ratio for $R=8,K=4$ is less than 0.19%. This is sufficient for production clusters.
@@ Appendix D Orchestration For Fat-Tree
Output: Placement scheme maximizing GPU utilization
@@
Fat-Tree topology is another common data center topology. A typical training strategy for this topology aims to maximize the bandwidth utilization under ToR (Top of Rack) Switches. Using Meta’s two-stage clos topology(Gangidi et al., 2024a) as a reference, it can be observed that there is an attempt to run CP under ToR.
Deployment Strategy: Assuming there are $p$ nodes under each ToR, nodes with the same index under each ToR are deployed along the same parallel sub-line, and the $p$ sub-lines are connected end-to-end, as shown in Figure 7. The training strategy involves running CP $p$ across the sub-lines and running TP within them.
@@
    Aggregation-Switches Domain Constraint: The coverage domian of a group of Aggregation Switches is limited, meaning that TP groups spanning across Aggregation Switches domains would result in cross-rail traffic, which should be avoided as much as possible.
@@
    TP Group Alignment Constraint: A CP Group consists of TP Groups across parallel sub-lines. To keep CP traffic within the ToR, the TP Groups must be aligned. If a node fails under one ToR, all nodes under that ToR are considered failed, expanding the failure radius by a factor of $p$.
@@ Appendix E Additional Simulation Results for Fault Resilience
This section presents additional simulation results related to §6.2. Figure 16 shows the variation of the GPU waste ratio over time under the production fault trace. Figure 17 presents the CDF data for the GPU waste ratio. Figure 18 illustrates the waste GPU ratio for different HBD architectures under various node failure rates, including the results for TP-8 to TP-64. Figure 19 shows the proportion of job-fault waiting time relative to total time for different job scales. All the aforementioned experiments include results for TP-8, TP-16, TP-32, and TP-64 configurations.
@@
Figure 16. GPU waste ratio over production fault trace, 4 GPU node.
@@
Figure 17. CDF of GPU waste ratio over production fault trace, 4 GPU node.
@@
Figure 18. GPU wastes ratio with different GPU fault ratio, 4-GPU node.
@@
Figure 19. Job fault-waiting duration with different levels of job-scale, 4 GPU node

### v4

@@ Appendix B GPT-MoE Architecture
Appendix C Theoretical analysis of GPU waste ratio for InfiniteHBD
The number of backup lines, given by $22K-$ 2, will significantly influence the fault tolerance of InfiniteHBD. We use the expectation of waste ratio caused by GPU failure and the fragmentation problem to evaluate this design, the result is shown in Table 7.
For a single working node in the middle of the line, the count of breakpoints $B$ on its two sides has the expectation as:
@@ Appendix C Theoretical analysis of GPU waste ratio for InfiniteHBD
Where $P_{s}$ is the failure probability of GPU node, and $\eta$ is the count of nodes.
@@
When the distance between two nodes is $\beta\geq K$, the breakpoints among them can be considered independent. Once the distance $\beta<K$, as all nodes in this range are connected to these two nodes, there will be no breakpoints between them. So, the expectation is less than two independent nodes. Then,
@@
For a LLM job which require a ring communication size (TP .etc) as $N_{t}$, InfiniteHBD will cut the whole line topology into several sub-lines with the length of $N_{t}/R$. Once InfiniteHBD is cutting a new sub-line from the remaining nodes in the line, all $N_{t}$ GPU will be wasted when one break point exist in the middle of this sub-line required, shown in Figure 19. Then the expectation for waste GPU caused by one single break point is:
@@
In our trace from a 3K-GPU cluster over 348 days, the 99th percentile (p99) failure rate for 8-GPU nodes is 7.22%. Therefore, we set the node failure rate $P_{s}=7.22\%$ when $R=8$. Based on the calculations in Appendix §A , given a GPU failure rate of $0.93\%$, we derive the corresponding node failure rate $P_{s}=3.67\%$ for $R=4$. If a TP32 job is running on InfiniteHBD (i.e., $N_{t}=32$), plugging these values into Equation (1), we obtain the upper bound of the expected waste ratio for different configurations, as shown in Table 7.
@@
R=4 $7.54\%$ $0.28\%$ $1.02\times 10^{-4}$
R=8 $25.02\%$ $1.81\%$ $0.13\%$
Table 7. Upper bound for waste ratio expectation of GPU, where GPU failure rate is 0.93% and $N_{t}$ is 32
For example, the 4-GPU node ($R=4$), 3-bundle ($K=3$) design incurs less than 0.28% additional GPU waste, while the waste ratio for the $R=8,K=4$ configuration remains below 0.13%. Both are sufficiently low for production cluster deployment.
@@ Appendix D Orchestration For Fat-Tree
Output: Placement scheme that maximizes GPU utilization
@@
The Fat-Tree is another common topology used in data centers. A typical training strategy for this topology aims to maximize the bandwidth utilization under ToR (Top of Rack) Switches. Using Meta’s two-stage clos topology(Gangidi et al., 2024) as a reference, it can be observed that there is an attempt to run CP under ToR Switches.
Deployment Strategy. Assuming there are $p$ nodes under each ToR, nodes with the same index under each ToR are deployed along the same parallel sub-line, and the $p$ sub-lines are connected end-to-end, as shown in Figure 7. The training strategy involves running CP $p$ across the sub-lines and running TP within them.
@@
    Aggregation-Switches Domain Constraint: The coverage domain of a group of Aggregation Switches is limited, meaning that TP groups spanning across Aggregation Switches domains would result in cross-rail traffic, which should be avoided as much as possible.
@@
    TP Group Alignment Constraint: A CP Group consists of TP Groups across parallel sub-lines. To keep CP traffic within the ToR switches, the TP Groups must be aligned. If a node fails under one ToR, all nodes under that ToR are considered to have failed, expanding the failure radius by a factor of $p$.
@@ Appendix E Additional Simulation Results for Fault Resilience
This section presents additional simulation results related to §6.2. Figure 20 shows the variation of the GPU waste ratio over time under the production fault trace. Figure 21 presents the CDF data for the GPU waste ratio. Figure 22 illustrates the GPU waste ratio for different HBD architectures under various node failure rates, including the results for TP-8 to TP-64. Figure 23 shows the proportion of job-fault waiting time relative to total time for different job scales. All the aforementioned experiments include results for TP-8, TP-16, TP-32, and TP-64 configurations.
@@
Figure 20. GPU waste ratio over production fault trace, 4-GPU node.
@@
Figure 21. CDF of GPU waste ratio over production fault trace, 4-GPU node.
@@
Figure 22. GPU waste ratio with different GPU fault ratio, 4-GPU node.
@@
Figure 23. Job fault-waiting duration under different job scales, 4-GPU node.
@@ Appendix F Detailed Cost and Power consumption Analysis
Appendix G AllToAll Communication of InfiniteHBD
Nowadays, training MoE models has become one of the primary application scenarios for GPU clusters. For large MoE models such as DeepSeek-v3 (DeepSeek-AI et al., 2025), both training and inference heavily rely on expert parallelism (EP). As analyzed earlier (§2.3), while TP can still effectively support MoE model training, for the sake of generality, we have theoretically explored how InfiniteHBD can be utilized to support EP AllToAll communication. Based on the OCSTrx Fast Switch Mechanism, we propose a method to enable the Binary Exchange AllToAll algorithm, though this approach has not yet been validated on real hardware.
G.1. OCSTrx Fast Switch Mechanism
In the original assumption, InfiniteHBD maintains a fixed topology configuration during any collective communication operation, meaning that only pre-activated links are used for communication. These links form a ring topology, which is inefficient for supporting AllToAll communication. This limitation is particularly common in traditional optical interconnect networks, where the end-to-end reconfiguration latency of OCS—composed of both hardware switching latency and control plane latency—is typically on the order of milliseconds, and in some cases even minutes (Urata et al., 2022), making it impractical to dynamically reconfigure for each communication pair.
However, OCSTrx adopts a more simplified hardware architecture that significantly reduces the reconfiguration latency compared to centralized OCS-based switches. Specifically, the hardware switching latency of our OCSTrx module is only 60–80 µs. Moreover, the control plane latency can be optimized and minimized by preloading Top-Session configurations in its control module and triggering flow switching as needed. As a result, the total end-to-end reconfiguration latency of OCSTrx is reduced to approximately 60-80 $\mu s$. We refer to this combined capability as the OCSTrx Fast Switch mechanism.
The Fast Switch mechanism allows us to fully utilize all backup links instead of being restricted to pre-activated ones. This means that active links can be dynamically configured in real-time based on any communication pair’s requirements, thereby improving the efficiency of AllToAll communication.
G.2. Binary Exchange AllToAll Algorithm
Fast Switch enables the system to fully utilize all available links. However, since nodes equipped with OCSTrx have a limited radix, the topology cannot achieve a fully connected structure at any arbitrary scale or location, but rather supports only sparse interconnects. Additionally, architectures like InfiniteHBD do not support node-level loopback at arbitrary scales.
Due to these limitations, AllToAll algorithms based on fully connected topologies (such as those used in NCCL (NVIDIA, 2020)) are not suitable for sparse interconnect environments. Similarly, the lack of node-level loopback makes algorithms that rely on sequential communication infeasible, even if they are inherently sparse. For instance, algorithms like Bruck (Bruck et al., 1997) and Pairwise Exchange (Thakur and Gropp, 2003), which depend on node-level loopback communication, cannot be directly applied to InfiniteHBD.
To better accommodate the interconnect characteristics of InfiniteHBD, we adopt the Binary Exchange AllToAll algorithm. In Binary Exchange, a node $i$ only communicates with node $i\oplus 2^{k}$. For example, in a communication group of 8 nodes, node 0 only exchanges data with nodes 1, 2, and 4 in different rounds. This communication pattern ensures sparsity and eliminates the need for node-level loopback, making it highly compatible with the topology of InfiniteHBD.
The detailed procedure of Binary Exchange is shown in Algorithm 6. Suppose the AllToAll group consists of $p$ nodes, and each node initially stores $p\cdot m$ units of data. Each node maintains the following two variables:
  • $Msg$: Stores all data currently held by the node, including both its original data and any received data.
  • $Commset$: Records the set of nodes that have already exchanged data (either directly or through intermediate nodes). Initially, this set contains only the node’s own index.
The algorithm runs for $\log_{2}p$ rounds. In the $k$-th step ($1\leq k\leq\log_{2}p$), node $i$ exchanges data with node $r=i\oplus 2^{\log_{2}p-k}$:
  1. (1)
    Node $i$ sends the following data fragment to node $r$:
    $Msg[Commset][r\mathbin{\&}2^{\log_{2}p-k}:r\mathbin{\&}2^{\log_{2}p-k}+2^{\log_{2}p-k}]$
    and receives data from node $r$. The transmitted data size per round is $\frac{p\cdot m}{2}$.
  2. (2)
    The communication set is updated:
    $i.Commset=i.Commset\cup r.Commset$
The total execution time of Binary Exchange is:
$T=\sum_{i=1}^{\log_{2}p}\left(t_{s}+t_{w}\frac{pm}{2}\right)=t_{s}\log_{2}p+\frac{1}{2}t_{w}mp\log_{2}p=O(p\log_{2}p)$
where $t_{s}$ is the transmission setup time and $t_{w}$ is the per-unit data transfer time.
This algorithm achieves a significantly lower communication complexity, making it well-suited for sparse interconnect architectures without node-level loopback, such as InfiniteHBD. Compared to the $O(p^{2})$ complexity of ring-AllToAll when Fast Switch is not used, Binary Exchange reduces the complexity to $O(p\log_{2}p)$, greatly improving communication efficiency.
Input: Group size $p$, Node index $i$, Initial message $m_{init}$
Output: AllToAll personalized data
Initialize $Msg[i]=m_{init}$, $Commset=\{i\}$;
for  $k$ in $1,...,log_{2}p$ do
       $r=i\oplus 2^{log_{2}p-k}$;
       $m_{send}=\{Msg[m][n]|m\in Commset,n\in[r\mathbin{\&}2^{\log_{2}p-k},r\mathbin{\&}2^{\log_{2}p-k}+2^{\log_{2}p-k})\}$;
       Node $i$ sends message $m_{send}$ to node $r$;
       Node $i$ receives message $m_{recv}$ from node $r$;
       Update $Msg$ with $m_{recv}$;
       $Commset=Commset\cup r.Commset$;
return $\{Msg[m][i]|m\in Commset\}$
Algorithm 6 Binary-Exchange-AllToAll
Figure 24. Illustration of Binary Exchange AllToAll algorithm under TP4+EP4 configuration and InfiniteHBD topology for AllToAll.
G.3. InfiniteHBD Topology for AllToAll
To support dynamically scalable Binary Exchange AllToAll, we propose a novel topology based on InfiniteHBD.
The topology retains the one-dimensional arrangement of nodes in InfiniteHBD but modifies the wiring pattern. Since in Binary Exchange AllToAll, node $i$ only communicates with node $i\oplus 2^{k}$, the distance between any two communicating nodes satisfies:
$\Delta=i-(i\oplus 2^{k})\in\{+2^{k},-2^{k}\}$
Based on this, instead of connecting each node to neighbors at distances ±$\pm 1,2,...,K$ as in the original K-Hop Ring topology, we rewire nodes to connect at distances ±$\pm 1,2,4,...,2^{K-1}$, aligning with the Binary Exchange AllToAll communication pattern.
This topology supports 2D parallelism of hybrid TP+EP in InfiniteHBD, where TP is the inner-layer parallelism and EP is the outer-layer parallelism. Figure 24 illustrates a TP4+EP4 configuration based on a 4-GPU Node setup, where TP4 is applied within each 4-GPU node, and EP4 is applied across four such nodes. The figure shows the steps of the Binary Exchange AllToAll algorithm used for EP4 communication: in Step 1, Node 0 exchanges data with Node 2, and Node 1 with Node 3; in Step 2, Node 0 exchanges data with Node 1, and Node 2 with Node 3.
However, this topology fundamentally emphasizes a trade-off between AllToAll performance and fault resilience, thereby introducing the following significant challenges:
  • Additional GPU forwarding overhead: Since interconnects are not GPU-level, some traffic must be forwarded through intermediate GPUs, introducing extra communication overhead.
  • Coupling between TP and EP: In 4-GPU Node, the number of OCSTrx bundles is limited, allowing each node to connect at most to ±$\pm 1,2,4,8$ distant nodes. As a result, the TP group size affects the EP group size, constrained by: $TP_{\text{size}}\times EP_{\text{size}}\leq 64$. This limits general applicability in some scenarios. The 8-GPU Node alleviates this issue, as it supports up to 8 OCSTrx bundles, enabling connections at distances ±$\pm 1,2,4,8,16,32,64,128$, relaxing the constraint to: $TP_{\text{size}}\times EP_{\text{size}}\leq 2048$, making it suitable for most scenarios.
  • High scheduling complexity: Node failures affect not only their own availability but also reduce bandwidth on backup links. Scheduling algorithms must simultaneously optimize GPU utilization and inter-node bandwidth, increasing system complexity.